A researcher reports an F-statistic (F-ratio) with df = (3, 18) from an independent-measures ANOVA. Based on the df values, how many treatments were compared in the study and what was the total number of subjects participating in the study? Group of answer choices 4 treatments and 21 subjects 3 treatments and 22 subjects 3 treatments and 21 subjects 4 treatments and 22 subjects
Answer #1

Based on df values

Treatment df = 3

Error df = 18

So Total df = Treatment df + Error df = 3+18 = 21

Therefore

Total number of treatments= Treatment df +1 = 3+1 = 4

Total number of observations = Total df +1 = 21 + 1 =22

Answer : 4 treatments and 22 subjects

​Please give thumbs up if you like the answer.